A function is defined by the polynomial 𝑓(đ‘„) = 3đ‘„ 4 − 4đ‘„ 3 − 12đ‘„ 2 + 8. Find and classify all the stationary points of f(x)


1
Expert's answer
2021-06-10T09:49:41-0400

Solution:-


"f(x)=3x^4-4x^3-12x^2+8"

Domain: "(-\infin, \infin)"



"f'(x)=(3x^4-4x^3-12x^2+8)'""=12x^3-12x^2-24x""=12x(x^2-x-2)"

Find the critical number(s)



"f'(x)=0=>12x(x^2-x-2)=0""12x(x+1)(x-2)=0"

Critical numbers: "-1,0,2."

If "x<-1, f'(x)<0, f(x)" decreases.

If "-1<x<0, f'(x)>0, f(x)" increases.

If "0<x<2, f'(x)<0, f(x)" decreases.

If "x>2, f'(x)>0, f(x)" increases.


"f(-1)=3"

"f(0)=8"

"f(2)=-24"

The function "f(x)" has a local maximum with value of "8" at "x=0."


The function "f(x)" has a local minimum with value of "3" at "x=-1."


The function "f(x)" has a local minimum with value of "-24" at "x=2."

"x=2."